OpenStudy (anonymous):

HELP! I need to simplify the complex fraction. Please Explain the problem so I can use this one as a example. 19/21/5/7

OpenStudy (anonymous):

(19/21)/(5/7)....when dividing fractions, " flip" what you are dividing by, then multiply 19/21 x 7/5 (you see how I flipped the number that I was dividing by. Now just multiply 19/21 x 7/5 = 19/15 = 1 4/15